OpenStudy (anonymous):

x^4 - 33x^2 - 108 = 0

OpenStudy (anonymous):

let \[u=x^2\]

OpenStudy (anonymous):

\[u^2-33u-108=0\]

OpenStudy (anonymous):

-108=-36*3 -36+3=-33 \[u^2+3u-36u+108\]

OpenStudy (anonymous):

\[u^2+3u-36u-108\]*

OpenStudy (anonymous):

\[u(u+3)-36(u+3)\]

OpenStudy (anonymous):

can you finish this?

OpenStudy (anonymous):

yes i got it sorry i forgot to close it.